# Reducing classical communication costs in multiplexed quantum repeaters using hardware-aware quasi-local policies

### v2

@@ Appendix C Entanglement distillation with the BBPSSW protocol
Now, when there are multiple (more than two) links that can be distilled, and we consider the pairwise BBPSSW protocol, the question of what is the optimal distillation policy naturally arises, i.e., what is the optimal pairing of links when performing entanglement distillation of more than two links. This question has been addressed in existing literature; see, e.g., Refs. [17, 84]. Here, we consider the following policies.
@@ C.1 The pumping policy
Figure 19: Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. C.2. Here we use the banded policy to perform distillation. The swap-distill policy is not useful beyond initial age $m^{\star}/2$, whereas distill-swap can extend this threshold to $m^{\star}$.
@@
Figure 20: Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. C.2. Now, instead of the banding policy, we use the pumping policy for distillation. As with banded purification, swap-distill is not useful beyond an initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold, but now not quite up to $\approx m^{\star}$. This is a consequence of the fact that the pumping policy has a strict limit on the fidelity that can be achieved with an increasing number of channels, as shown in Eq. (55).
All of the analysis above was done considering a idealised situation in which links of the same age are available always for distillation. This is of course not necessarily true in practice. In practical settings, links will have to be used as and when they are produced. Therefore, it is important to look also at the previously mentioned pumping policy.
@@
Appendix D Distill-swap vs. swap-distill policy: analytical results
Here we analytically compare the performance of the distill-swap and swap-distill policies in the case of three nodes. To simplify the analysis, we restrict ourselves to the cases with two and four channels, and consider all elementary links to be active simultaneously with the same fidelity. The general result for an arbitrary number of channels is difficult to calculate analytically given the large number of possible trajectories that can lead to the successful generation of at least one end-to-end link. Seeing in Appendix C.2 that the banded policy performs better than pumping, in this section we look exclusively at the banded policy. The performance of the policies is evaluated by the success probability of establishing at least one end-to-end entangled link and by the average fidelity of an end-to-end link.
D.1 Two-channel case
Let $f_{0}$ be the initial fidelity of both the active links on each side given in terms of the initial age $m_{0}$ according to Eq. (25) as $f_{0}=f(m_{0})=\frac{1}{4}(1+3\mathrm{e}^{-m_{0}/m^{\star}})$.
@@ D.1 Two-channel case
Next, we consider the swap-distill policy. For this policy, two cases may arise depending on the number of successful entanglement swaps. Let $p_{sd}^{(1)}$, $f_{sd}^{(1)}$ and $p_{sd}^{(2)}$, $f_{sd}^{(2)}$ be the success probability and fidelity, respectively, of the two cases. The two cases are based on whether one or both of the entanglement swaps succeed. Then, the total probability of producing at least one end-to-end link is the sum of two cases and is given by
@@
D.2 Four-channel case
Now, for the four-channel case, let us start with the distill-swap policy. Recall that under the banding policy, at most three rounds of distillation can be performed with four links. Based on this, there are three trajectories can lead to one end-to-end link, and the corresponding success probabilities and fidelities are derived below. We combine them to obtain the total probability and weighted (expected) fidelity of an end-to-end link, in a manner analogous to (61) and (62). Throughout these derivations, we let $f_{0}$ be the initial fidelity of all links, we let $p_{d}\equiv P_{\text{distill}}(f_{0},f_{0})$ and $f_{d}\equiv P_{\text{distill}}(f_{0},f_{0})$ be the success probability and fidelity, respectively, after the first round of distillation. Similarly, we let $p_{d_{2}}\equiv P_{\text{distill}}(f_{d},f_{d})$ and $f_{d_{2}}\equiv F_{\text{distill}}(f_{d},f_{d})$ be the success probability and fidelity, respectively, after the second round of distillation. The corresponding ages of the links after each round are denoted by $m_{d}\equiv f^{-1}(f_{d})$ and $m_{d_{2}}\equiv f^{-1}(f_{d_{2}})$, where we recall the function $f(m)=\frac{1}{4}(1+3\exp(-m/m^{\star}))$ and its inverse $f^{-1}(F)=\lceil-m^{\star}\log((4F-1)/3)\rceil$.
@@ D.2 Four-channel case
    Two entanglement swapping attempts succeed, and then one distillation attempt is made and succeeds. There are now $\binom{4}{2}=6$ possible ways to obtain two successful entanglement swaps, which means the the success probability $p_{sd}^{(2)}$ and fidelity $f_{sd}^{(2)}$ in the case are given by
@@
    Thus, the success probability and fidelity is a sum (and weighted sum) of these two sub-cases, and is given by:
@@
    Case 4: All 4 entanglement swapping attempts succeed. Again two sub-cases arise—one where all three subsequent distillation attempts succeed, and second where the one of the two first round of distillation attempts fail (can happen in two ways), and there is no scope to distill further. Thus, the success probability and fidelity is a sum (and weighted sum) of these two sub-cases, and is given by:
@@
Figure 21: Comparison of the distill-swap and swap-distill policies for a three-node chain with four channels connecting the neighboring nodes, as described in Sec. D. We look at the probability of successfully producing at least one end-to-end entangled link (left) and the average age of that link (when successfully produced) (right). For distillation, we use the banded policy. We see again that the swap-distill policy is not useful beyond and initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold up to $m^{\star}$ by using an increasing number of channels. $p_{sw}=50\%$ for these plots.
In Fig. 21, we show the results for four channels. It is clear from Fig. 21 that swap-distill has a higher probability of success compared to distill-swap when the initial age of the links is less than $m^{\star}/2$, beyond which it fails to succeed. Unlike distillation, success probability of entanglement swapping is independent of the quality of the used links. This is the primary reason for an asymmetry in the success rates of swap-distill and distill-swap policies. The probability of at least one entanglement swap being successful increases exponentially with the number of attempts available, whereas the probability of all distillations being successful falls with an increasing number of attempts, irrespective of the ages of the links involved. Thus, it is clear that if the aim is to generate at least one end-to-end link as quickly as possible, i.e., if the aim is to minimize waiting time, the swap-distill is the policy of choice. At the same time, distill-swap succeeds until a larger initial age threshold; thus, if the elementary links are of low fidelity to begin with, distilling first might be judicious. Also, in terms of fidelity enhancement (age reduction), distill-swap performs better. These conclusions are also consistent with the results in Fig. 19 and 20.
Appendix E Other policies for multiplexing based on swap-asap
@@
Figure 22: (Top) Performance of mixed-weight swap-asap multiplexing policies in terms of average waiting time and age of an end-to-end link, as a function of the mixing parameter $a$. FN ($a=1$) is nearly optimal for waiting time reduction, but a nearly equal mixture of SN and FN ($a=0.4$) seems to minimize the average age of an end-to-end link. (Bottom) The figures of merit for the random-priority swap-asap policy show a much more monotonic trend. FN ($r=1$) is optimal for waiting time and SN ($r=0$) is optimal for average age.
@@
Here, we also want to consider the question of optimality of SN and FN swap-asap in terms of reducing the average age and waiting time for an end-to-end link. Of course, a full optimization could be done by explicitly defining the figure of merit as a cost function and then using tools such reinforcement learning or linear programs to find the optimal entanglement distribution policy within our MDP framework. As was shown in previous works like Refs. [45, 56], the number of MDP states grows exponentially with the number of links and also with the cutoff. Thus, finding the optimal policy explicitly in the multiplexing case is quite challenging, and we leave that as an aim for future work. Here, we ask a limited question. Let us consider a policy that intends to match links for entanglement swapping in such a way that the total length of the virtual links (if entanglement swaps were to be successful) is maximum. We define this policy as FN optimal, or FN-OPT policy. In other words, FN-OPT tries all pairings of links for entanglement swapping, calculates the sum of the lengths of the anticipated virtual links, and then chooses the set of pairings that maximizes this sum. Such a pairing could be different from the pairings provided by the usual FN policy, in which the longest link on each side is paired with one another, followed by the second longest, and so on. For example, if we have two links of ages $7$ and $3$ on each side on a node, then two entanglement swaps can be performed in total. Let us say that the cutoff is 10. The standard FN pairing would be $(7,7)$ and $(3,3)$. However, this would mean that the first entanglement swap will never be attempted, because we know that it will fail (as described earlier, the swapped link has age $m_{1}+m_{2}$, and we only perform swaps if $m_{1}+m_{2}<m^{\star}$). Therefore, the sum of lengths of anticipated virtual links in this case will be $3+3=6$. FN-OPT, on the other hand, would choose the pairings $(7,3)$ and $(7,3)$, because the sum of anticipated lengths would be $10+10=20$. The intuition for considering FN-OPT is the same as the standard FN policy, i.e., prioritizing forming long links reduces the average waiting time. In a similar way, an SN-OPT policy can be defined, which would minimize the sum of the ages of the anticipated links.
@@
Figure 23: We compare the performance of FN to FN-OPT and SN to SN-OPT in terms of average waiting time and average age of an end-to-end link. The much simpler and standard FN and SN policies are nearly optimal in both cases.
@@
Appendix F Classical communication costs for local, quasi-local and global policies
In this appendix, we aim to discuss some of the details related to our model for calculating the classical communications costs for different policies. As was mentioned in Sec. IV, based on the amount of information available to nodes when making policy decisions, policies can be classified as local, quasi-local, and global. We remind the reader that in our model, for all policies, no CC is allowed within one MDP step. Consequently, if at the beginning of an MDP step it is determined that multiple interdependent swaps or distillation attempts are to be made, the failure of one leads to the failure of all dependent swaps or distillation attempts.
Figure 24: (Left) A generic network evolution trajectory. The CC cost is estimated in our model as the length of the longest link involved in a swap in a given time-step. The swap at node 6 leads to a CC cost of 3 units, where 1 unit is equal to the heralding time for the elementary link. (Right) Evolution of a section of the chain with no connections initially. The end nodes stop waiting for CC if there is a gap in the stream of CC packets they receive.
Consider the example of a generic network state as shown in Fig. 24 (left). We ignore distillation in this discussion, the CC cost for which is much easier to estimate and has been explained in Sec. IV. Let us focus on the actions of node 6 and isolate one of the channels between the nodes of the chain. Node 6 knows at the beginning of the MDP step that it is connected to node 3. Now it will try to generate an elementary link with node 7. As soon as such a link is generated, it will perform a swap. Some of the time evolution trajectories that can emerge are as follows.
To elaborate on such a scheme further, we give another example in Fig. 24 (right). We consider a portion of the network in which there are no links to begin with and show how a connected segment of the chain is established, with each node aware of its connections.
At the beginning of the protocol, each node tries to generate an elementary link. Suppose all of these attempts succeed. Now, each node will attempt swaps simultaneously, because the information about their respective elementary link generation success is available in one time step (heralding time). Now, if one of the swaps fails all of them must fail, and this information will be available at each node, allowing them to restart their memories in the next time step. This will also mark the end of one round of the protocol for this portion of the chain and no extra CC will be required. The other possibility is that all the swaps have succeeded, in which case the internal nodes will be free to restart without the need for any CC. They will just send their measurement outcomes to their next node, that is the node they were connected to at the beginning of the outcome, this CC signal will now be relayed along the chain with more information being added at each node until all this information reaches both the end nodes (see Fig. 24 (right)). One of the end nodes can keep utilizing this classical information to perform the local operations necessary to create an end-to-end entangled link. (Which one? We can have a set convention that the node on the left, i.e., one receiving CC signals from the right of it, will do Pauli rotations.) As more and more CC has reached them, they get informed about the success and failure of swaps that occurred farther and farther away, and they get to know that they now share a link with some far-off node that has a certain age. The end nodes will stop waiting for more CC once there is a gap of one time step in the CC packets reaching them. This will indicate that they have received all CC from the connected region of the chain.

### v3

@@ Acknowledgements.
When there are multiple (more than two) links that can be distilled, and we consider the pairwise BBPSSW protocol, the question of what is the optimal distillation policy naturally arises, i.e., what is the optimal pairing of links when performing entanglement distillation of more than two links. This question has been addressed in existing literature; see, e.g., Refs. [17, 88]. Here, we consider the following policies.
@@
Figure 18: Swap-then-distill versus distill-then-swap policy using banded approach to distillation. Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Quasi-local multiplexing policies of main text. Here we use the banded policy to perform distillation. The swap-distill policy is not useful beyond initial age $m^{\star}/2$, whereas distill-swap can extend this threshold to $m^{\star}$.
@@
Figure 19: Swap-then-distill versus distill-then-swap policy using pumping approach to distillation.Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described above. Now, instead of the banding policy, we use the pumping policy for distillation. As with banded purification, swap-distill is not useful beyond an initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold, but now not quite up to $\approx m^{\star}$. This is a consequence of the fact that the pumping policy has a strict limit on the fidelity that can be achieved with an increasing number of channels, as shown in Eq. (57).
All of the analysis above was done considering an idealized situation in which links of the same age are available always for distillation. This is of course not necessarily true in practice. In practical settings, links will have to be used as and when they are produced. Therefore, it is important to look also at the previously mentioned pumping policy.
@@
Supplementary Note 3: Distill-swap vs. swap-distill policy: analytical results
Here we analytically compare the performance of the distill-swap and swap-distill policies in the case of three nodes. To simplify the analysis, we restrict ourselves to the cases with two and four channels, and consider all elementary links to be active simultaneously with the same fidelity. The general result for an arbitrary number of channels is difficult to calculate analytically given the large number of possible trajectories that can lead to the successful generation of at least one end-to-end link. Seeing in Supplementary Note 2 that the banded policy performs better than pumping, in this section we look exclusively at the banded policy. The performance of the policies is evaluated by the success probability of establishing at least one end-to-end entangled link and by the average fidelity of an end-to-end link.
Two-channel case
Let $f_{0}$ be the initial fidelity of both the active links on each side given in terms of the initial age $m_{0}$ as $f_{0}=f(m_{0})=\frac{1}{4}(1+3\mathrm{e}^{-m_{0}/m^{\star}})$.
@@
Next, we consider the swap-distill policy. For this policy, two cases may arise depending on the number of successful entanglement swaps. Let $p_{sd}^{(1)}$, $f_{sd}^{(1)}$ and $p_{sd}^{(2)}$, $f_{sd}^{(2)}$ be the success probability and fidelity, respectively, of the two cases. The two cases are based on whether one or both of the entanglement swaps succeed. Then, the total probability of producing at least one end-to-end link is the sum of the two cases and is given by
@@
Four-channel case
Now, for the four-channel case, let us start with the distill-swap policy. Recall that under the banding policy, at most three rounds of distillation can be performed with four links. Based on this, there are three trajectories that can lead to one end-to-end link, and the corresponding success probabilities and fidelities are derived below. We combine them to obtain the total probability and weighted (expected) fidelity of an end-to-end link, in a manner analogous to (63) and (64). Throughout these derivations, we let $f_{0}$ be the initial fidelity of all links, we let $p_{d}\equiv P_{\text{distill}}(f_{0},f_{0})$ and $f_{d}\equiv P_{\text{distill}}(f_{0},f_{0})$ be the success probability and fidelity, respectively, after the first round of distillation. Similarly, we let $p_{d_{2}}\equiv P_{\text{distill}}(f_{d},f_{d})$ and $f_{d_{2}}\equiv F_{\text{distill}}(f_{d},f_{d})$ be the success probability and fidelity, respectively, after the second round of distillation. The corresponding ages of the links after each round are denoted by $m_{d}\equiv f^{-1}(f_{d})$ and $m_{d_{2}}\equiv f^{-1}(f_{d_{2}})$, where we recall the function $f(m)=\frac{1}{4}(1+3\exp(-m/m^{\star}))$ and its inverse $f^{-1}(F)=\lceil-m^{\star}\log((4F-1)/3)\rceil$.
@@
    Two entanglement swapping attempts succeed, and then one distillation attempt is made and succeeds. There are now $\binom{4}{2}=6$ possible ways to obtain two successful entanglement swaps, which means the success probability $p_{sd}^{(2)}$ and fidelity $f_{sd}^{(2)}$ in the case are given by
@@
    Thus, the success probability and fidelity are a sum (and weighted sum) of these two sub-cases, and are given by:
@@
    Case 4: All 4 entanglement swapping attempts succeed. Again, two sub-cases arise—one where all three subsequent distillation attempts succeed, and second where one of the two first rounds of distillation attempts fails (can happen in two ways), and there is no scope to distill further. Thus, the success probability and fidelity is a sum (and weighted sum) of these two sub-cases, and is given by:
@@
Figure 20: Analytical results for distill-swap and swap-distill policies. Comparison of the distill-swap and swap-distill policies for a three-node chain with four channels connecting the neighboring nodes, as described in Supplementary Note 3. We look at the probability of successfully producing at least one end-to-end entangled link (left) and the average age of that link (when successfully produced) (right). For distillation, we use the banded policy. We see again that the swap-distill policy is not useful beyond and initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold up to $m^{\star}$ by using an increasing number of channels. $p_{sw}=50\%$ for these plots.
In Fig. 20, we show the results for four channels. It is clear from Fig. 20 that swap-distill has a higher probability of success compared to distill-swap when the initial age of the links is less than $m^{\star}/2$, beyond which it fails to succeed. Unlike distillation, success probability of entanglement swapping is independent of the quality of the used links. This is the primary reason for an asymmetry in the success rates of swap-distill and distill-swap policies. The probability of at least one entanglement swap being successful increases exponentially with the number of attempts available, whereas the probability of all distillations being successful falls with an increasing number of attempts, irrespective of the ages of the links involved. Thus, it is clear that if the aim is to generate at least one end-to-end link as quickly as possible, i.e., if the aim is to minimize waiting time, the swap-distill is the policy of choice. At the same time, distill-swap succeeds until a larger initial age threshold; thus, if the elementary links are of low fidelity to begin with, distilling first might be judicious. Also, in terms of fidelity enhancement (age reduction), distill-swap performs better. These conclusions are also consistent with the results in Fig. 18 and 19.
Supplementary Note 4: Other policies for multiplexing based on swap-asap
@@
Figure 21: Performance of other multiplexing policies. (Top) Performance of mixed-weight swap-asap multiplexing policies in terms of average waiting time and age of an end-to-end link, as a function of the mixing parameter $a$. FN ($a=1$) is nearly optimal for waiting time reduction, but a nearly equal mixture of SN and FN ($a=0.4$) seems to minimize the average age of an end-to-end link. (Bottom) The figures of merit for the random-priority swap-asap policy show a much more monotonic trend. FN ($r=1$) is optimal for waiting time and SN ($r=0$) is optimal for average age.
@@
Here, we also want to consider the question of optimality of SN and FN swap-asap in terms of reducing the average age and waiting time for an end-to-end link. Of course, a full optimization could be done by explicitly defining the figure of merit as a cost function and then using tools such as reinforcement learning or linear programs to find the optimal entanglement distribution policy within our MDP framework. As was shown in previous works like Refs. [45, 56], the number of MDP states grows exponentially with the number of links and also with the cutoff. Thus, finding the optimal policy explicitly in the multiplexing case is quite challenging, and we leave that as an aim for future work. Here, we ask a limited question. Let us consider a policy that intends to match links for entanglement swapping in such a way that the total length of the virtual links (if entanglement swaps were to be successful) is maximum. We define this policy as FN optimal, or FN-OPT policy. In other words, FN-OPT tries all pairings of links for entanglement swapping, calculates the sum of the lengths of the anticipated virtual links, and then chooses the set of pairings that maximizes this sum. Such a pairing could be different from the pairings provided by the usual FN policy, in which the longest link on each side is paired with one another, followed by the second longest, and so on. For example, if we have two links of ages $7$ and $3$ on each side on a node, then two entanglement swaps can be performed in total. Let us say that the cutoff is 10. The standard FN pairing would be $(7,7)$ and $(3,3)$. However, this would mean that the first entanglement swap will never be attempted, because we know that it will fail (as described earlier, the swapped link has age $m_{1}+m_{2}$, and we only perform swaps if $m_{1}+m_{2}<m^{\star}$). Therefore, the sum of lengths of anticipated virtual links in this case will be $3+3=6$. FN-OPT, on the other hand, would choose the pairings $(7,3)$ and $(7,3)$, because the sum of anticipated lengths would be $10+10=20$. The intuition for considering FN-OPT is the same as the standard FN policy, i.e., prioritizing forming long links reduces the average waiting time. In a similar way, an SN-OPT policy can be defined, which would minimize the sum of the ages of the anticipated links.
@@
Figure 22: Performance of our multiplexing policies compared to their optimal versions. We compare the performance of FN to FN-OPT and SN to SN-OPT in terms of average waiting time and average age of an end-to-end link. The much simpler and standard FN and SN policies are nearly optimal in both cases.